# Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

### v2

@@ 4 Results
Table 5: Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions. Effect sizes are shown in Figure˜5.
@@
Effect Size Generally modest, consistent with recent large-scale findings [48]
@@
Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions. Table˜5 gives a summary of our key results. Figure˜5 complements this with a depiction of effect sizes for the controlled factors.
Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility. The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}\leq 0.001$). See also Table˜5.
4.1 H1: Phish Scale Effect
Figure 6: Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention.
H1 was supported. Figure˜6 illustrates the comparison of effect sizes across different factors in phishing susceptibility. The NIST Phish Scale successfully predicted user behavior. Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p << 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity. Higher difficulty emails consistently produced higher click rates across all experimental conditions:
  • Easy difficulty: 7.0% average click rate
  • Medium difficulty: 8.7% average click rate
  • Hard difficulty: 15.0% average click rate
This statistically significant difficulty gradient confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts.
4.2 H2: Overall Training Effects
H2 was not supported. Training did not produce statistically significant improvements in either click reduction or reporting behavior. Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, pp=0.450) or report rates (F(2, 12086) = 0.874, pp=0.417). This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [48].
4.3 H3: Interactive Training Effects
@@
This finding contrasts with some prior work suggesting benefits of interactive training components [19], but aligns with recent large-scale studies showing limited training effectiveness regardless of modality [48].
4.4 H4: Training-Difficulty Interactions
@@
Coefficient analysis revealed some interesting interaction patterns:
  • Training Only × Hard: -9.67 percentage points (pp=0.002)
  • Training + Exercise × Hard: -9.09 percentage points (pp=0.004)
These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant. It is worth considering that the training vendor created the phishing templates and training materials, which may introduce alignment between the cues emphasized in training and those present in the simulated emails. This alignment could artificially enhance the apparent effectiveness of training for certain email types, particularly those closely modeled after the vendor’s examples. As a result, the observed reductions in click rates for hard emails may partially reflect this overlap rather than a generalizable improvement in user detection skills. Future studies should consider using independently developed phishing templates or varying template sources to better assess the robustness of training effects across diverse phishing scenarios.
4.5 H5: Reporting Timeliness
H5 was partially supported, though with important caveats about sample size limitations. Our analysis of reporting timeliness revealed mixed patterns in organizational-level temporal protection across treatment groups. Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment, indicating generally rapid user response to suspicious emails.
At the organizational level, we observed notable differences in inoculation patterns—campaigns where reports preceded clicks, enabling organizational threat mitigation before exploitation. Training Only achieved the highest inoculation rate at 54.5% of campaigns, with a median organizational response time of -0.33 hours (reports preceding clicks). In contrast, the Control group achieved inoculation in 42.9% of campaigns with a median response time of 8.09 hours, while Training + Exercise showed 36.4% inoculation with a median response time of 1.21 hours.
However, these differences were not statistically significant due to necessarily small sample sizes at the campaign level (n=33 campaigns total). The Organizational Inoculation Index revealed substantial organizational resilience independent of individual training effectiveness, with over one-third of all template-group combinations demonstrating proactive reporting patterns. While Training Only showed numerically superior temporal performance, the limited campaign-level sample size prevents definitive conclusions about training effects on organizational-level reporting timeliness.
@@ 4.5 H5: Reporting Timeliness
These findings provide the first empirical measurement of organizational temporal security dynamics, directly addressing Steves et al.’s research question about whether "time to first report is sooner than time to first click" [88]. The results suggest that organizational-level protective behaviors may operate independently of individual training effectiveness, though larger-scale studies are needed to establish statistical significance of these temporal patterns.
4.6 Other Measurements
The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness. In our study, subjects received the phishing simulation email between 0–4 months after the training. Figure˜7 shows the click rate by treatment group, discriminating by lure difficulty. We observe no discernible trend in the results as a function of time. Given the result of the visual inspection, we did not perform statistical tests of this aspect.
Figure 7: Temporal Impact on Detection. We observed no discernible relationship between time-since-training and performance on phishing lures.
4.7 Effect Sizes and Practical Significance
@@
While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01 for all main effects), raising questions about practical significance [90]. The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
Training effects, where present, were modest in magnitude. Even statistically significant interaction effects represented changes of less than 10 percentage points, suggesting that while training may provide some benefits, these are limited in scope and may not justify the substantial organizational investment required for comprehensive training programs [48].
@@
These findings underscore the need for realistic expectations about training effectiveness and suggest that organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection against phishing threats [37].
5 Discussion
Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches. While we observed some statistically significant effects, their practical significance raises important questions about the return on investment for organizational security awareness programs.
5.1 The Reality of Training (In)Effectiveness and Organizational Security
@@
Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417). The small effect sizes we observed ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact. This aligns with recent large-scale evidence from Ho et al. [48] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
Our findings suggest that many organizations may be operating under a dangerous illusion of training effectiveness. The cybersecurity community must confront this reality by demanding more rigorous evaluation standards and refusing to accept vendor claims that cannot be substantiated under realistic threat conditions. Only by acknowledging the true limitations of current training approaches can organizations make informed decisions about resource allocation and security strategy [37].
The Possibility of Training Metrics Manipulation: Our findings suggest the possibility that organizations may inadvertently—or deliberately—manipulate phishing training metrics to achieve desired outcomes. The strong predictive power of the NIST Phish Scale (F(2, 12086) = 41.415, p < 0.001) demonstrates that practitioners can significantly influence training results simply by adjusting the difficulty of phishing lures used in simulations.
Our observation informs broader concerns about “compliance theater” in cybersecurity [44], where organizations optimize for metrics that satisfy regulatory requirements rather than actual security improvement. Recent work by Ho et al. [48] and Lain et al. [53, 54] provides additional evidence that training effectiveness is often overstated. Our findings demonstrate a specific mechanism through which such overstatement may occur: the systematic use of phishing lures that fail to represent realistic attack sophistication.
Such selection may not be the fault of the organizations alone. We note that many vendors of phishing training allow organizations to select from libraries of phishing templates without standardized difficulty assessment. This enables the selection—accidental or deliberate—of easier lures, resulting in inflated success metrics [18].
5.2 Recommendations for Phish Resilience
Layered Defense Strategy: Our findings add to the evidence that organizational phishing defense strategy must treat training as one of many layers of defense. The strong effect of phishing difficulty on user susceptibility (F(2,12086) = 41.415, p < 0.001) indicates that attacks are likely to succeed regardless of training efforts.
Evidence-Based Implementation Framework: Organizations seeking meaningful security assessment must resist the temptation to achieve favorable metrics through lure selection bias. Our results support the following three-part framework:
(1) Technical controls first. As humans cannot be trained to reliably detect phishing, automated defenses are in order. Prior work recommends investing primarily in filtering, attachment analysis, URL protection, and multi-factor or phishing-resistant authentication [4].
(2) Targeted training. Our results suggest that generic company-wide training is ineffective [48]. Prefer focused, role- or scenario-specific training where prior evidence suggests impact [82, 90]. Use standardized difficulty frameworks (e.g., the NIST Phish Scale) for evaluation.
(3) Continuous assessment. Regularly compare training costs and outcomes against alternative investments using standardized metrics that reflect behavior and resource efficiency.
In phishing training assessment, guided by our methods, we recommend that organizations:
  • Use standardized difficulty frameworks like the NIST Phish Scale to ensure representative threat simulation
  • Report results separately by difficulty level rather than aggregating across lure types
  • Benchmark against external standards rather than setting arbitrary internal targets
  • Focus on improvement trends over time rather than absolute performance thresholds
We do not recommend eliminating phishing training, but rather to right-size its role within a defense-in-depth strategy guided by evidence rather than marketing claims.
The Evolving Threat Landscape: Our findings must be interpreted within the context of rapidly evolving phishing threats. The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete [4, 80, 9]. When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it may become unreasonable to expect users to reliably identify deceptive content. [^5] One of the authors recently failed his organization’s annual phishing training, which was experimenting with GenAI-based lures.
The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks. Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [30]. Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load [33, 32]. Training represents only a single layer of defense against these multifaceted human vulnerabilities.
This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity operations community must reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity. Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms to reduce reliance on the “human firewall”.
Implications for Theory and Practice: Our findings have significant implications for both cybersecurity theory and organizational practice. The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats. The modest effect sizes ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
The interaction patterns we identified—where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails—highlight the complex and sometimes counterintuitive effects of security awareness interventions. These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
For practitioners, our results support a realistic reassessment of training investments within broader security strategies. Organizations should view awareness programs as one component of layered defense rather than primary protection mechanisms. The substantial resources typically devoted to comprehensive training initiatives might achieve better risk reduction when reallocated toward technical controls, incident response capabilities, or more targeted interventions [37].
5.3 Future Work
Our findings open several important avenues for future research that could advance understanding of phishing training effectiveness and organizational cybersecurity resilience.
AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks that may lack the flaws traditionally used to assess difficulty, such as poor grammar or implausible scenarios [80]. This calls into question the continued relevance of the NIST Phish Scale, which assumes such imperfections are present. Although some phishing campaigns will continue to use low-budget approaches with human-induced errors, future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts and whether new paradigms are needed to simulate and defend against these advanced threats.
Automated Difficulty Assessment: The manual process of rating phishing lures using the NIST Phish Scale is subject to rater subjectivity and does not scale well to large deployments. Future work should explore automated or semi-automated systems—potentially powered by machine learning—to assess lure difficulty by evaluating features such as cue visibility and premise plausibility. Additionally, comparing lure success rates in randomized control groups with sufficiently large populations could offer a data-driven alternative for estimating difficulty. These approaches could improve consistency, scalability, and real-world applicability of phishing difficulty assessments.
Alternative Training Modalities and Interventions: While our study focused on lecture-based and interactive training approaches commonly offered by vendors, other intervention types remain understudied. Future research should investigate the effectiveness of just-in-time training, personalized adaptive tools that adjust to users’ individual risk profiles [82], and hybrid models that combine human training with technical controls. Emerging interventions—such as AI-driven, real-time reviews of links and emails—also warrant exploration. Even widely deployed tools like spelling and grammar checkers may offer untapped potential as passive defenses against phishing.

### v3

@@ 4. Results
Table 4. Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
@@
Effect Size Consistent with large-scale findings (hoUnderstandingEfficacyPhishing2025, )
  Training Effect $\eta^{2}=0.000$ (negligible)
  Difficulty Effect $\eta^{2}=0.007$ (meaningful)
  Training $\times$ Difficulty $\eta^{2}=0.001$ (negligible)
@@
Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions. Table 4 gives a summary of our key results. Appendix 4 complements this with a depiction of effect sizes for the controlled factors.
4.1. H1: Phish Scale Effect
Figure 2. Comparison of effect sizes of phishing susceptibility. Bars show the variance explained by phishing difficulty (per the NIST Phish Scale) for each training intervention. :
H1 was supported. Figure 2 illustrates the comparison of effect sizes across different factors in phishing susceptibility. The NIST Phish Scale successfully predicted user behavior. Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086)=41.415, p << 0.001, $\eta^{2}$=0.007), validating the scale’s utility for measuring phishing complexity. Higher difficulty emails consistently produced higher click rates across all experimental conditions; Easy difficulty: 7.0% average click rate, Medium difficulty: 8.7% average click rate, Hard difficulty: 15.0% average click rate.
4.2. H2: Overall Training Effects
H2 was not supported. Training did not produce statistically significant improvements in either click reduction or reporting behavior. Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, pp=0.450) or report rates (F(2, 12086) = 0.874, pp=0.417). This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches (hoUnderstandingEfficacyPhishing2025, ).
4.3. H3: Interactive Training Effects
@@
This finding contrasts with some prior work suggesting benefits of interactive training components (carellaImpactSecurityAwareness2017, ), but aligns with recent large-scale studies showing limited overall training effectiveness (hoUnderstandingEfficacyPhishing2025, ).
4.4. H4: Training-Difficulty Interactions
@@
Coefficient analysis revealed some interesting interaction patterns: Training Only × Hard: -9.67 percentage points (pp=0.002); Training + Exercise × Hard: -9.09 percentage points (pp=0.004) These significant negative interactions suggest that training may provide some protection against the most difficult phishing emails, even though overall training effects were not significant. Higher alignment messages with fewer cues are the most deceptive (butaviciusBreachingHumanFirewall2015, ).
4.5. H5: Reporting Timeliness
H5 was partially supported, though with important caveats about sample size limitations. Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment. We analyzed 17 templates that received both clicks and reports. Overall, 52.9% achieved inoculation (reports preceded clicks), with median Organizational Inoculation Index (OII) of +0.4 hours. By treatment group, Control achieved inoculation on 66.7% of templates (median OII: +0.3 hours), Training Only on 75% (median OII: +1.37 hours), and Training + Exercise on only 28.6% (median OII: -0.2 hours). These differences were not statistically significant due to limited sample size. For successfully inoculated templates, the median OII was approximately 30 minutes, suggesting automated response mechanisms would be necessary to mitigate threats before exploitation.
@@ 4.5. H5: Reporting Timeliness
These findings provide the first empirical measurement of organizational temporal security dynamics, responding to Steves et al.’s research question about whether “time to first report is sooner than time to first click” (stevesCategorizingHumanPhishing2020, ). The results suggest that organizational-level protective behaviors operate independently of individual training effectiveness. Studying the opportunities and limitations of this feedback loop is a promising direction for further study.
4.6. Effect Sizes and Practical Significance
@@
While some coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01), raising questions about practical significance (sumnerExaminingFactorsImpacting2022, ). The only meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity. Even statistically significant interaction effects represented changes of less than 10 percentage points.
@@
These findings underscore the need for realistic expectations about training effectiveness and suggest organizations should complement training with other defense mechanisms rather than relying on awareness programs as primary protection (franzSoK, ).
5. Discussion
5.1. Implications for Practice
@@
After deploying regulation-compliant training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417). The small effect sizes we observed ($\eta^{2}$ ¡ 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact. This result aligns with recent large-scale evidence from Ho et al. (hoUnderstandingEfficacyPhishing2025, ) and challenges the optimistic assessments promoted by cybersecurity training vendors.
The predictive power of the NIST Phish Scale (F(2, 12086)=41.415, p ¡ 0.001) demonstrates that practitioners can significantly influence training results by adjusting the difficulty of phishing lures used in simulations. Our data shows click rates vary dramatically based on lure difficulty: from 7.0% for easy emails to 15.0% for hard emails. Organizations setting specific performance targets (such as click-through rates around 2%) may achieve these by using low-difficulty lures that fail to represent sophisticated attacks organizations actually face (alkhalilPhishingAttacksRecent2021, ). This creates a misalignment between perceived and actual security posture.
These findings suggest several actionable opportunities for both training vendors and organizational security teams. Vendors can enhance their value by incorporating standardized measures of lure difficulty and adaptive testing frameworks aligned with models such as the NIST Phish Scale. This would produce interpretable results that organizations could use to calibrate difficulty levels to track progress over time, personalize training to employee risk profiles, and benchmark outcomes across departments or industries. However, the relatively low impact of phishing training suggests that awareness programs must be part of a layered defense strategy, complemented by technical controls and incident response preparedness (franzSoK, ), to gain risk reduction rather than nominal compliance (haneyComplianceImpactTracing2024, ). In this way, security awareness initiatives can evolve from static, one-size-fits-all modules into data-driven, context-aware components of an organization’s continuous improvement process.
5.2. Future Work